OpenStudy (anonymous):

A group of friends were voting for their favorite song: two-fifths chose song A, one-sixth of the remaining friends chose song B, and the rest chose song C. If there were thirty friends in all, how many chose song C? ten twelve fifteen eighteen

OpenStudy (imstuck):

You need to figure out first how many friends chose song A. Then subtract that number from the total number to get the remaining number of friends. Let's start there. Can you figure out how many friends chose song A?

OpenStudy (anonymous):

4

OpenStudy (imstuck):

No, I'm sorry that's not correct. If there are 30 total people and 2/5 chose song A, how many people chose song A? Can you come up with an expression for that?

OpenStudy (anonymous):

is it 8 people

OpenStudy (tkhunny):

30(2/5) = 12 that chose A. Do the arithmetic. Stop guessing.

OpenStudy (anonymous):

I dont know the arithmetic im stupid :/

OpenStudy (imstuck):

Ok, here's your set up for the number of people who chose song A: \[\frac{ 2 }{ 5}*30\]You an reduce between the 5 and the 30 and get 2 * 6 = 12. 12 friends chose song A. If 12 chose song A, how many friends were left to choose B or C?

OpenStudy (imstuck):

Please don't say you're stupid. That is not true at all. Being stupid would be not understanding it and then not caring.

OpenStudy (anonymous):

I think its B

OpenStudy (imstuck):

No, B is not correct. 12 chose A. That leaves 30 - 12 = 18 friends. They tell you that 1/6 of the remaining 18 chose B. 1/6 * 18 = 3. The first 12 chose A then 3 chose B, so that accounts for 15 friends. All the rest of the 30 chose C. If 15 chose A or B, then how many chose C?

OpenStudy (imstuck):

Can you get it?

OpenStudy (anonymous):

is the answer 12

OpenStudy (imstuck):

No, it's 15. That is choice C or 3 down from the top.

OpenStudy (anonymous):

thank you even though I couldnt really figure it out :(

OpenStudy (imstuck):

If you have another one, post it here. Practice makes perfect! And you're welcome!

OpenStudy (anonymous):

I have one more A piece of music is made up of m measures. If each measure has 3 beats, how many total beats are in the piece of music? three m three plus m m divided by three m